OpenStudy (anonymous):

what is the solution of this equation? 1w=-17

sammixboo (sammixboo):

Just divide 1 on both sides :) \(1w~=~-17\) \(\frac{1w}{1}~=~\frac{-17}{1}\) \(w~=~\frac{-17}{1}\) So what is -17/1?

OpenStudy (anonymous):

chosenmatt PLEASE HELP ME WITH MY QUESTION

OpenStudy (anonymous):

w=-17

OpenStudy (anonymous):

-17/1=-17